Suppose that 62% of people like chocolate ice cream. If you pick two people at random, what is the probability that they both like chocolate ice cream? Give your answer as a decimal (to at least 3 places) or fraction.
Answer #1

Here, n = 2, p = 0.62, (1 - p) = 0.38 and x = 2
As per binomial distribution formula P(X = x) = nCx * p^x * (1 - p)^(n - x)

We need to calculate P(X = 2)
P(X = 2) = 2C2 * 0.62^2 * 0.38^0
P(X = 2) = 0.384